OpenStudy (anonymous):

Let y = (4x+(8/x))^-4 . Find derivative

OpenStudy (goformit100):

This question can be done in two methods

OpenStudy (goformit100):

You can use the chain rule for differentiation, or you can show the steps

OpenStudy (yttrium):

\[y' = -4 (4x+ \frac{ 8 }{ x })^{-5} (4-\frac{ -8 }{ x^2 })\]

OpenStudy (yttrium):

Do you know to simplify it now?

OpenStudy (anonymous):

I figured out the first part but how did you get: (4−(-8/x^2)) ?

OpenStudy (anonymous):

it doesnt tell me to simplify

OpenStudy (yttrium):

It's the derivative of the term inside the exponent. It also needs to be derived.

OpenStudy (anonymous):

ok

OpenStudy (goformit100):

Madam do you know the chain rule of differentiation ?

OpenStudy (anonymous):

is that the same as the constant multiple rule?

OpenStudy (goformit100):

it's a Method which find the derivative at one go.